\begin{thebibliography}{41}
\providecommand{\natexlab}[1]{#1}
\providecommand{\url}[1]{\texttt{#1}}
\expandafter\ifx\csname urlstyle\endcsname\relax
  \providecommand{\doi}[1]{doi: #1}\else
  \providecommand{\doi}{doi: \begingroup \urlstyle{rm}\Url}\fi

\bibitem[Ankan and Panda(2015)]{ankan2015pgmpy}
Ankur Ankan and Abinash Panda.
\newblock pgmpy: Probabilistic graphical models using {P}ython.
\newblock In \emph{Proceedings of the 14th Python in Science Conference}, pages 6--11, 2015.

\bibitem[Atienza et~al.(2022)Atienza, Bielza, and Larra{\~n}aga]{atienza2022semiparametric}
David Atienza, Concha Bielza, and Pedro Larra{\~n}aga.
\newblock Semiparametric {B}ayesian networks.
\newblock \emph{Information Sciences}, 584:\penalty0 564--582, 2022.

\bibitem[Barons et~al.(2022)Barons, Mascaro, and Hanea]{barons2022balancing}
Martine~J Barons, Steven Mascaro, and Anca~M Hanea.
\newblock Balancing the elicitation burden and the richness of expert input when quantifying discrete {B}ayesian networks.
\newblock \emph{Risk Analysis}, 42\penalty0 (6):\penalty0 1196--1234, 2022.

\bibitem[Bodewes and Scutari(2021)]{bodewes2021learning}
Tjebbe Bodewes and Marco Scutari.
\newblock Learning {B}ayesian networks from incomplete data with the node-average likelihood.
\newblock \emph{International Journal of Approximate Reasoning}, 138:\penalty0 145--160, 2021.

\bibitem[Constantinou et~al.(2021)Constantinou, Liu, Chobtham, Guo, and Kitson]{constantinou2021large}
Anthony~C Constantinou, Yang Liu, Kiattikun Chobtham, Zhigao Guo, and Neville~K Kitson.
\newblock Large-scale empirical validation of {B}ayesian network structure learning algorithms with noisy data.
\newblock \emph{International Journal of Approximate Reasoning}, 131:\penalty0 151--188, 2021.

\bibitem[Constantinou et~al.(2023)Constantinou, Guo, and Kitson]{constantinou2023impact}
Anthony~C Constantinou, Zhigao Guo, and Neville~K Kitson.
\newblock The impact of prior knowledge on causal structure learning.
\newblock \emph{Knowledge and Information Systems}, 65\penalty0 (8):\penalty0 3385--3434, 2023.

\bibitem[de~Beaufort et~al.(2015)de~Beaufort, Sedki, and Fontenelle]{de2015inference}
Louis~B de~Beaufort, Karima Sedki, and Guy Fontenelle.
\newblock Inference reasoning on fishers' knowledge using {B}ayesian causal maps.
\newblock \emph{Ecological Informatics}, 30:\penalty0 345--355, 2015.

\bibitem[Ducamp et~al.(2020)Ducamp, Gonzales, and Wuillemin]{ducamp2020agrum}
Gaspard Ducamp, Christophe Gonzales, and Pierre-Henri Wuillemin.
\newblock a{GrUM/pyAgrum}: A toolbox to build models and algorithms for probabilistic graphical models in {P}ython.
\newblock In \emph{International Conference on Probabilistic Graphical Models}. PMLR, 2020.

\bibitem[Epskamp et~al.(2012)Epskamp, Cramer, Waldorp, Schmittmann, and Borsboom]{epskamp2012qgraph}
Sacha Epskamp, Ang{\'e}lique~OJ Cramer, Lourens~J Waldorp, Verena~D Schmittmann, and Denny Borsboom.
\newblock qgraph: Network visualizations of relationships in psychometric data.
\newblock \emph{Journal of Statistical Software}, 48:\penalty0 1--18, 2012.

\bibitem[Geiger and Heckerman(1994)]{geiger1994learning}
Dan Geiger and David Heckerman.
\newblock Learning {G}aussian networks.
\newblock In \emph{Proceedings of the 10th Conference on Uncertainty in Artificial Intelligence}, pages 235--243, 1994.

\bibitem[Hanea et~al.(2015)Hanea, Napoles, and Ababei]{hanea2015non}
Anca Hanea, Oswaldo~Morales Napoles, and Dan Ababei.
\newblock Non-parametric {B}ayesian networks: Improving theory and reviewing applications.
\newblock \emph{Reliability Engineering \& System Safety}, 144:\penalty0 265--284, 2015.

\bibitem[Heckerman and Geiger(1995)]{heckerman1995learning1}
David Heckerman and Dan Geiger.
\newblock Learning bayesian networks: A unification for discrete and {G}aussian domains.
\newblock In \emph{Proceedings of 11th Conference on Uncertainty in Artificial Intelligence}, pages 274--284, 1995.

\bibitem[Heckerman et~al.(1995)Heckerman, Geiger, and Chickering]{heckerman1995learning}
David Heckerman, Dan Geiger, and David~M Chickering.
\newblock Learning {B}ayesian networks: The combination of knowledge and statistical data.
\newblock \emph{Machine Learning}, 20:\penalty0 197--243, 1995.

\bibitem[H{\o}jsgaard(2012)]{hojsgaard2012graphical}
S{\o}ren H{\o}jsgaard.
\newblock {Graphical independence networks with the gRain package for R}.
\newblock \emph{Journal of Statistical Software}, 46:\penalty0 1--26, 2012.

\bibitem[Hosseini and Ivanov(2020)]{hosseini2020bayesian}
Seyedmohsen Hosseini and Dmitry Ivanov.
\newblock Bayesian networks for supply chain risk, resilience and ripple effect analysis: A literature review.
\newblock \emph{Expert Systems with Applications}, 161:\penalty0 113649, 2020.

\bibitem[Johnson et~al.(2014)Johnson, Abal, Ahern, and Hamilton]{jon}
Sandra Johnson, Eva Abal, Kathleen Ahern, and Grant Hamilton.
\newblock From science to management: Using {B}ayesian networks to learn about {L}yngbya.
\newblock \emph{Statistical Science}, 29\penalty0 (1):\penalty0 36--41, 2014.

\bibitem[Kabir and Papadopoulos(2019)]{kabir2019applications}
Sohag Kabir and Yiannis Papadopoulos.
\newblock Applications of {B}ayesian networks and {P}etri nets in safety, reliability, and risk assessments: A review.
\newblock \emph{Safety science}, 115:\penalty0 154--175, 2019.

\bibitem[Kaikkonen et~al.(2021)Kaikkonen, Parviainen, Rahikainen, Uusitalo, and Lehikoinen]{kaikkonen2021bayesian}
Laura Kaikkonen, Tuuli Parviainen, Mika Rahikainen, Laura Uusitalo, and Annukka Lehikoinen.
\newblock Bayesian networks in environmental risk assessment: A review.
\newblock \emph{Integrated Environmental Assessment and Management}, 17\penalty0 (1):\penalty0 62--78, 2021.

\bibitem[Kelly et~al.(2013)Kelly, Jakeman, Barreteau, Borsuk, ElSawah, Hamilton, Henriksen, Kuikka, Maier, Rizzoli, van Delden, and Voinov]{kelly2013selecting}
Rebecca~A Kelly, Anthony~J Jakeman, Olivier Barreteau, Mark~E Borsuk, Sondoss ElSawah, Serena~H Hamilton, Hans~J{\o}rgen Henriksen, Sakari Kuikka, Holger~R Maier, Andrea~Emilio Rizzoli, Hedwig van Delden, and Alexey~A Voinov.
\newblock Selecting among five common modelling approaches for integrated environmental assessment and management.
\newblock \emph{Environmental Modelling \& Software}, 47:\penalty0 159--181, 2013.

\bibitem[Kitson et~al.(2023)Kitson, Constantinou, Guo, Liu, and Chobtham]{kitson2023survey}
Neville~K Kitson, Anthony~C Constantinou, Zhigao Guo, Yang Liu, and Kiattikun Chobtham.
\newblock A survey of {B}ayesian network structure learning.
\newblock \emph{Artificial Intelligence Review}, 56\penalty0 (8):\penalty0 8721--8814, 2023.

\bibitem[Koller and Friedman(2009)]{koller2009}
Daphne Koller and Nir Friedman.
\newblock \emph{Probabilistic graphical models: Principles and techniques}.
\newblock MIT press, 2009.

\bibitem[Kratzer et~al.(2023)Kratzer, Lewis, Comin, Pittavino, and Furrer]{kratzer2023additive}
Gilles Kratzer, Fraser Lewis, Arianna Comin, Marta Pittavino, and Reinhard Furrer.
\newblock Additive {B}ayesian network modeling with the {R} package abn.
\newblock \emph{Journal of Statistical Software}, 105:\penalty0 1--41, 2023.

\bibitem[Kuipers et~al.(2022)Kuipers, Suter, and Moffa]{kuipers2022efficient}
Jack Kuipers, Polina Suter, and Giusi Moffa.
\newblock Efficient sampling and structure learning of {B}ayesian networks.
\newblock \emph{Journal of Computational and Graphical Statistics}, 31\penalty0 (3):\penalty0 639--650, 2022.

\bibitem[Kyrimi et~al.(2021)Kyrimi, McLachlan, Dube, Neves, Fahmi, and Fenton]{kyrimi2021comprehensive}
Evangelia Kyrimi, Scott McLachlan, Kudakwashe Dube, Mariana~R Neves, Ali Fahmi, and Norman Fenton.
\newblock A comprehensive scoping review of {B}ayesian networks in healthcare: Past, present and future.
\newblock \emph{Artificial Intelligence in Medicine}, 117:\penalty0 102108, 2021.

\bibitem[Leonelli et~al.(2020)Leonelli, Riccomagno, and Smith]{leonelli2020coherent}
Manuele Leonelli, Eva Riccomagno, and Jim~Q Smith.
\newblock Coherent combination of probabilistic outputs for group decision making: An algebraic approach.
\newblock \emph{OR Spectrum}, 42\penalty0 (2):\penalty0 499--528, 2020.

\bibitem[Leonelli et~al.(2023)Leonelli, Ramanathan, and Wilkerson]{leonelli2023sensitivity}
Manuele Leonelli, Ramsiya Ramanathan, and Rachel~L Wilkerson.
\newblock Sensitivity and robustness analysis in {B}ayesian networks with the bnmonitor {R} package.
\newblock \emph{Knowledge-Based Systems}, 278:\penalty0 110882, 2023.

\bibitem[Lin et~al.(2020)Lin, Neil, and Fenton]{lin2020improved}
Peng Lin, Martin Neil, and Norman Fenton.
\newblock Improved high dimensional discrete {B}ayesian network inference using triplet region construction.
\newblock \emph{Journal of Artificial Intelligence Research}, 69:\penalty0 231--295, 2020.

\bibitem[Liu and Constantinou(2022)]{liu2022greedy}
Yang Liu and Anthony~C Constantinou.
\newblock Greedy structure learning from data that contain systematic missing values.
\newblock \emph{Machine Learning}, 111\penalty0 (10):\penalty0 3867--3896, 2022.

\bibitem[Marcot and Penman(2019)]{marcot2019advances}
Bruce~G Marcot and Trent~D Penman.
\newblock Advances in {B}ayesian network modelling: Integration of modelling technologies.
\newblock \emph{Environmental modelling \& software}, 111:\penalty0 386--393, 2019.

\bibitem[Moe et~al.(2021)Moe, Carriger, and Glendell]{moe2021increased}
Jannicke~S Moe, John~F Carriger, and Miriam Glendell.
\newblock Increased use of {B}ayesian network models has improved environmental risk assessments.
\newblock \emph{Integrated Environmental Assessment and Management}, 17\penalty0 (1):\penalty0 53--61, 2021.

\bibitem[Nodelman et~al.(2002)Nodelman, Shelton, and Koller]{nodelman2002continuous}
Uri Nodelman, Christian~R Shelton, and Daphne Koller.
\newblock Continuous time {B}ayesian networks.
\newblock In \emph{Proceedings of the 18th Conference on Uncertainty in Artificial Intelligence}, pages 378--387, 2002.

\bibitem[Nyberg et~al.(2022)Nyberg, Nicholson, Korb, Wybrow, Zukerman, Mascaro, Thakur, Oshni~Alvandi, Riley, Pearson, Morris, Herrmann, Azad, Bolger, Hahn, and Lagnado]{nyberg2022bard}
Erik~P Nyberg, Ann~E Nicholson, Kevin~B Korb, Michael Wybrow, Ingrid Zukerman, Steven Mascaro, Shreshth Thakur, Abraham Oshni~Alvandi, Jeff Riley, Ross Pearson, Shane Morris, Matthieu Herrmann, A~K~M Azad, Fergus Bolger, Ulrike Hahn, and David Lagnado.
\newblock {BARD: A structured technique for group elicitation of Bayesian networks to support analytic reasoning}.
\newblock \emph{Risk Analysis}, 42\penalty0 (6):\penalty0 1155--1178, 2022.

\bibitem[Pearl(1988)]{pearl1988probabilistic}
Judea Pearl.
\newblock Probabilistic reasoning in intelligent systems: Network of plausible inference.
\newblock \emph{Morgan Kaufmann}, 1988.

\bibitem[P{\'e}rez-Bernab{\'e} et~al.(2016)P{\'e}rez-Bernab{\'e}, Fern{\'a}ndez, Rum{\'\i}, and Salmer{\'o}n]{perez2016parameter}
Inmaculada P{\'e}rez-Bernab{\'e}, Antonio Fern{\'a}ndez, Rafael Rum{\'\i}, and Antonio Salmer{\'o}n.
\newblock Parameter learning in hybrid {B}ayesian networks using prior knowledge.
\newblock \emph{Data Mining and Knowledge Discovery}, 30:\penalty0 576--604, 2016.

\bibitem[Peters et~al.(2017)Peters, Janzing, and Sch\"olkopf]{Peters2017}
Jonas Peters, Dominik Janzing, and Bernhard Sch\"olkopf.
\newblock \emph{Elements of causal inference: Foundations and learning algorithms}.
\newblock MIT Press, 2017.

\bibitem[Pitchforth and Mengersen(2013)]{pitchforth2013proposed}
Jegar Pitchforth and Kerrie Mengersen.
\newblock A proposed validation framework for expert elicited {B}ayesian networks.
\newblock \emph{Expert Systems with Applications}, 40\penalty0 (1):\penalty0 162--167, 2013.

\bibitem[Rudin(2019)]{rudin2019}
Cynthia Rudin.
\newblock Stop explaining black box machine learning models for high stakes decisions and use interpretable models instead.
\newblock \emph{Nature Machine Intelligence}, 1\penalty0 (5):\penalty0 206--215, 2019.

\bibitem[Scutari(2010)]{scutari2010}
Marco Scutari.
\newblock Learning {B}ayesian networks with the {bnlearn} {R} package.
\newblock \emph{Journal of Statistical Software}, 35\penalty0 (3):\penalty0 1--22, 2010.

\bibitem[Scutari et~al.(2019)Scutari, Graafland, and Guti{\'e}rrez]{scutari2019learns}
Marco Scutari, Catharina~E Graafland, and Jos{\'e}~M Guti{\'e}rrez.
\newblock Who learns better {B}ayesian network structures: Accuracy and speed of structure learning algorithms.
\newblock \emph{International Journal of Approximate Reasoning}, 115:\penalty0 235--253, 2019.

\bibitem[Shiguihara et~al.(2021)Shiguihara, Lopes, and Mauricio]{shiguihara2021dynamic}
Pedro Shiguihara, Alneu De~Andrade Lopes, and David Mauricio.
\newblock Dynamic {B}ayesian network modeling, learning, and inference: A survey.
\newblock \emph{IEEE Access}, 9:\penalty0 117639--117648, 2021.

\bibitem[Yu et~al.(2020)Yu, Moharil, and Blair]{yu2020bayesnetbp}
Han Yu, Janhavi Moharil, and Rachael~Hageman Blair.
\newblock {BayesNetBP: An R package for probabilistic reasoning in Bayesian networks}.
\newblock \emph{Journal of Statistical Software}, 94:\penalty0 1--31, 2020.

\end{thebibliography}

\documentclass[twoside,11pt]{article}

\usepackage{blindtext}
\usepackage{bm}
\usepackage{amsmath}
\usepackage{jmlr2e}

\usepackage{lastpage}
\ShortHeadings{bnRep: A Repository of Bayesian Networks}{Manuele Leonelli}
\firstpageno{1}

\begin{document}

\title{bnRep: A Repository of Bayesian Networks from the Academic Literature}

\author{\name Manuele Leonelli\email manuele.leonelli@ie.edu \\
       \addr School of Science and Technology\\
       IE University\\
       Madrid, Spain}

\editor{}

\maketitle

\begin{abstract}
Bayesian networks (BNs) are widely used for modeling complex systems with uncertainty, yet repositories of pre-built BNs remain limited. This paper introduces \texttt{bnRep}, an open-source R package offering a comprehensive collection of documented BNs, facilitating benchmarking, replicability, and education. With over 200 networks from academic publications, \texttt{bnRep} integrates seamlessly with \texttt{bnlearn} and other R packages, providing users with interactive tools for network exploration.
\end{abstract}

\begin{keywords}
Bayesian networks, Probabilistic graphical models, R package, Repository
\end{keywords}

\section{Introduction}
Bayesian networks (BNs) \citep{pearl1988probabilistic} are a powerful machine learning model widely used in practice due to their ability to model complex systems with uncertainty and interpretability. Their widespread application spans diverse fields such as engineering reliability \citep{kabir2019applications}, environmental sciences \citep{kaikkonen2021bayesian}, healthcare \citep{kyrimi2021comprehensive}, and supply chain management \citep{hosseini2020bayesian}, among others. They are implemented in various software tools, including R packages (e.g. \texttt{bnlearn} \citealp{scutari2010}; \texttt{gRain} \citealp{hojsgaard2012graphical}), Python libraries (e.g. \texttt{pgmpy} \citealp{ankan2015pgmpy}; \texttt{pyAgrum} \citealp{ducamp2020agrum}), and standalone software (e.g. AgenaRisk, BayesiaLab, BayesServer, Hugin Expert, GeNie Modeler, and Netica)\footnote{Available at \url{https://www.agena.ai}, \url{https://www.bayesia.com}, \url{https://www.bayesserver.com}, \url{https://www.hugin.com}, \url{https://www.bayesfusion.com/genie/}, and \url{https://www.norsys.com/netica.html}, respectively.}, making them accessible to a wide range of users and applications. Despite their widespread application, BNs remain an active research area with ongoing advancements in learning algorithms \citep[e.g.][]{kuipers2022efficient,liu2022greedy}, inference techniques \citep[e.g.][]{lin2020improved}, and new types of networks \citep[e.g.][]{atienza2022semiparametric}.

Key reasons for the success of BNs include their explainability, flexibility, and status as the gold standard for causal modeling. Unlike black-box models, BNs offer an intuitive, transparent structure, aligning with trends in eXplainable AI (XAI) \citep{rudin2019}. They provide insights into cause-and-effect relationships rather than just correlations \citep{Peters2017}. BNs are also accessible to a broad audience, not requiring deep expertise in computer science or mathematics \citep{kelly2013selecting,moe2021increased}. Their modular design integrates data, expert knowledge, and model outputs \citep{leonelli2020coherent,marcot2019advances} and supports scenario and ``what-if" analyses for simulating interventions and predicting outcomes \citep[e.g.][]{pitchforth2013proposed}.

Despite the widespread use, flexibility, and numerous advantages of BNs, repositories of pre-built and documented BNs remain limited compared to other areas of machine learning. In fields like deep learning, there are extensive repositories such as TensorFlow Hub and PyTorch Hub that provide access to a wide range of pre-trained models\footnote{Available at \url{https://www.tensorflow.org/hub}and \url{https://pytorch.org/hub/}, respectively.}. Similarly, \texttt{scikit-learn} offers datasets and model repositories for classic machine learning algorithms, facilitating replication, benchmarking, and educational purposes. In contrast, repositories of BNs are relatively scarce, with some notable exceptions like the Bayesian Network Repository (including 31 BNs), the BNMA BN repository (including 75 BNs), and the BayesFusion Interactive Model Repository (including 46 BNs).\footnote{Available at \url{https://www.bnlearn.com/bnrepository/}, \url{https://bnma.co/bnrepo/}, and \url{https://repo.bayesfusion.com/}, respectively. } However, these collections are often smaller, provide less detailed documentation, and are not consistently updated to reflect recent advancements in BN research.

This gap motivated the creation of \texttt{bnRep}, an open-source R package that offers a comprehensive collection of documented BNs, enabling users to explore, compare, and apply them across various domains. The repository includes over 200 BNs from more than 150 different academic publications, each accompanied by detailed documentation. In the following sections, we provide a brief overview of BNs and present the key features and purposes of the \texttt{bnRep} package, highlighting why R was chosen as the development platform and how the package serves both research and practical needs.

\section{Bayesian networks}

While a full account of Bayesian networks can be found in several monographs \citep[e.g.][]{koller2009}, a brief overview is provided here to highlight the key concepts.

A BN gives a graphical representation of the relationship between a vector of variables of interest $\bm{Y}=(Y_1,\dots,Y_p)$ using a directed acyclic graph (DAG) $G$ and a factorization of the overall probability distribution $P(\bm{Y})$ in terms of simpler conditional distributions $P(Y_i\,|\, \bm{Y}_{\Pi_i})$, where $\bm{Y}_{\Pi_i}$ denotes the parents of $Y_i$ in $G$. More formally, the overall factorization of the probability distribution induced by the BN can be written as:
\begin{equation}
\label{eq:eq}
P_G(\bm{Y})=\prod_{i=1}^p P(Y_i\, |\,\bm{Y}_{\Pi_i}).
\end{equation}
This factorization has several advantages: it reduces the number of parameters to estimate, allows for easier expert elicitation of local relationships, and makes better use of available data by focusing only on the relevant variables and their parents.

The academic literature has mostly focused on three types of distributional assumptions depending on the nature of the variables: discrete, Gaussian, and conditional linear Gaussian
\citep[see e.g.][]{bodewes2021learning}. \textit{Discrete BNs} \citep{heckerman1995learning} are such that $\bm{Y}$ is a Multinomial random variable and the local distributions are defined as
\begin{equation}
\label{eq:1}
Y_i\,|\,\bm{Y}_{\Pi_i}\sim \text{Multi}(\theta_{ijk}), \mbox{ where } \theta_{ijk} = P(Y_i=j\,|\,\bm{Y}_{\Pi_i}=k).
\end{equation}
The parameters $\theta_{ijk}$ are usually reported in conditional probability tables, reporting the conditional probabilities for each parents' configuration.

\textit{Gaussian BNs} \citep{geiger1994learning} assume $\bm{Y}$ follows a multivariate Normal distribution and the local distributions are defined as linear regressions over the parents:
\begin{equation}
\label{eq:2}
Y_i\,|\, \bm{Y}_{\Pi_i} \sim \mathcal{N}\left(\beta_{i0} +\sum_{j\in\Pi_i}\beta_{ij}y_j, \sigma^2_i\right),
\end{equation}
where the $\beta$'s are the regression parameters and $\sigma^2_i>0$.

\textit{Conditional Linear Gaussian BNs} \citep{heckerman1995learning1} combine discrete and continuous random variables. Discrete $Y_i$'s can only have discrete parents and their distribution is as in Equation (\ref{eq:1}). Continuous $Y_i$'s can have both discrete and continuous parents, and their distribution is a mixture of Gaussian distributions (Equation \ref{eq:2}), one for each discrete parent configuration.

BNs can be constructed in various ways: they may be learned from data using algorithms that infer both the structure (DAG) and the conditional probabilities \citep{kitson2023survey,scutari2019learns}, elicited from expert knowledge \citep{barons2022balancing,nyberg2022bard}, or developed through a combination of the two. In practice, many models combine data-driven learning for the probabilistic parameters with expert-elicited structures or relationships, allowing for more accurate and interpretable representations of complex systems \citep{constantinou2023impact}.

\section{The bnRep package}

\subsection{Why R?}
R was chosen as the platform for \texttt{bnRep} due to its robust ecosystem for statistical modeling and seamless integration with existing tools for BNs. The \texttt{bnlearn} package \citep{scutari2010}, one of the most widely used packages for learning and inference in BNs, provides the core class objects for \texttt{bnRep}, directly supporting the three main types of BNs previously discussed. Additionally, \texttt{bnlearn} includes various exporting functions (such as \texttt{write.bif} and \texttt{write.net}), enabling users to easily transfer models between R and both Python and basically all standalone BN software. This ensures that models from \texttt{bnRep} can be utilized across different platforms and software environments, making the repository versatile for diverse applications and users' preferences.

The choice of R ensures that \texttt{bnRep} benefits from the CRAN ecosystem, which enforces clear documentation standards and regular package updates. This guarantees that \texttt{bnRep} is user-friendly, thoroughly documented, and accessible to a wide range of users. Moreover, \texttt{bnRep} is compatible with other R packages such as \texttt{gRain} \citep{hojsgaard2012graphical} and \texttt{BayesNetBP} \citep{yu2020bayesnetbp} for inference, \texttt{bnmonitor} \citep{leonelli2023sensitivity} for model diagnostics, and \texttt{qgraph} \citep{epskamp2012qgraph} for visualization, allowing for extended functionality within the R environment.

\subsection{Purposes of bnRep}

With growing interest in comparing the performance of structural learning algorithms \citep{constantinou2021large,scutari2019learns}, \texttt{bnRep} plays a central role in providing a benchmarked repository of models that enhances replicability and promotes collaboration in BN research. Researchers can test new algorithms against established models, ensuring consistent evaluation across studies. Additionally, the package serves as an educational tool, offering a diverse collection of models that help students and practitioners explore and learn about BNs across various domains \citep{de2015inference,jon}. By sharing knowledge through an accessible, well-documented, and up-to-date repository, \texttt{bnRep} fosters both innovation and cross-disciplinary applications of BNs.

\subsection{bnRep at a glance}

The \texttt{bnRep} package currently includes 214 BNs, with the number of networks constantly evolving as new models are added. Most networks come from papers published from 2020 onwards, and all are stored as \texttt{bn.fit} objects from the \texttt{bnlearn} package, which supports three types of BNs: discrete, Gaussian, and conditional linear Gaussian. These are the most widely used types of BNs in research and applications, making \texttt{bnlearn}'s support for these models crucial for \texttt{bnRep}'s extensive and versatile repository.

The package also includes the \texttt{bnRep\_summary} dataframe, which provides detailed information about each network. This includes the type of network, summaries of the DAG structure (form, number of nodes, edges, etc.), how the probabilities and DAG were defined (from data, expert knowledge, etc.), and the area of application (e.g., environmental science, engineering, medicine). Summaries of the repository and the networks' characteristics, along with visualizations, are available on the \texttt{bnRep} GitHub page (\url{https://github.com/manueleleonelli/bnRep}).

To facilitate interactive exploration, \texttt{bnRep} includes a Shiny app, accessible via the \texttt{bnRep\_app()} function. The app allows users to filter and explore the networks' database. It is freely available online at \url{https://manueleleonelli.shinyapps.io/bnRep/}, providing a user-friendly interface to access the repository without needing to install the R package.

\section{Conclusions and future directions}

The \texttt{bnRep} package provides a comprehensive, well-documented repository of BNs, making it a valuable resource for researchers and practitioners. By facilitating benchmarking, replicability, and education, \texttt{bnRep} supports the advancement of BN research and cross-field collaboration. Its integration with \texttt{bnlearn} ensures compatibility with widely-used tools, while the Shiny app enhances accessibility for users.

Future updates will expand the repository to include new types of networks, such as copula \citep{hanea2015non}, additive \citep{kratzer2023additive}, hybrid \citep[e.g.][]{perez2016parameter}, dynamic \citep[e.g.][]{shiguihara2021dynamic}, and continuous time BNs \citep{nodelman2002continuous}. The project’s open-source nature encourages contributions through GitHub, fostering community-driven growth. Additionally, benchmarking functions will be introduced, promoting the development of consistent evaluation methods in the BN field. In conclusion, \texttt{bnRep} not only addresses the current lack of documented BNs but also lays the foundation for establishing benchmarks and fostering innovation in BN research.

% Acknowledgements and Disclosure of Funding should go at the end, before appendices and references

% Manual newpage inserted to improve layout of sample file - not
% needed in general before appendices/bibliography.

\vskip 0.2in
\bibliography{sample}

\end{document}